Suppose rocket traveler Mary has a clock made on Earth. Every year on her birthday she sends a light signal to brother John on Earth. At what rate does John receive the signals during Mary’s outward journey


1
Expert's answer
2021-05-06T18:47:27-0400

The rate is "\frac{1}{\sqrt{1-\frac{v^2}{c^2}}}," where "v" is the Mary's speed.